### v2

@@ 4.4 Ablation Study
Settings. We conduct ablation studies of the tiny-size VRWKV on ImageNet-1K [8] to validate the effectiveness of various key components like Q-Shift and bidirectional attention. The experimental settings are consistent with Sec. 4.1.
Token Shift. We compare the performance of not using token shift, using the original shift method in RWKV [35], and our Q-Shift. As shown in Tab. 5, the variation in the shift method shows performance differences. Variant 1 without token shift leads to a poor performance of 71.5, which is 3.6 points lower than our model. Even with the use of our global attention, the model using the original token shift still has a 0.7-point gap with our model.
Bidirectional Attention. The bidirectional attention mechanism enables the model to achieve global attention while the original RWKV attention has a causal mask internally. The result of Variant 3 shows that the global attention mechanism brings a 2.3 points increase in the top-1 accuracy.
@@
Table 5: Ablation on key components of the proposed VRWKV. All models are trained from scratch on ImageNet-1K. The “original” means the original token shift in RWKV [35] which mixes the token in a single direction is used.
Effective Receptive Field (ERF). We analyze the impact of different designs on the ERF of models based on [11] and visualize it in Fig. 3(a). We visualize the ERF of the central pixel with an input size of 1024 × 1024. In  Fig. 3(a), “No Shift” represents the absence of the token shift method (Q-Shift), and “RWKV Attn” indicates using the original RWKV attention mechanism without our modifications for vision tasks. From the comparison in the figure, all models except the “RWKV Attn” one achieved global attention while the global capacity of the VRWKV-T model is better than that of ViT-T. Despite the assistance of Q-Shift, the central pixel in “RWKV Attn” still cannot attend to the pixels on the bottom of the image due to the large input resolution. The results of “No Shift” and Q-Shift show that the Q-Shift method expands the core range of the receptive field, enhancing the inductive bias of global attention.
Efficiency Analysis. We gradually increase the input resolution from $224\times 224$ to $2048\times 2048$ and compare the inference and memory efficiency of VRWKV-T and ViT-T. The results were tested on an Nvidia A100 GPU, as shown in Fig. 1. From the curves presented in Fig. 1(b), it is observed that at lower resolutions, such as $224\times 224$ with around 200 image tokens, VRWKV-T and ViT-T exhibit comparable memory usage, though VRWKV-T has a slightly lower FPS compared to ViT-T. However, with increasing resolution, VRWKV-T’s FPS rapidly exceeds that of ViT-T, thanks to its linear attention mechanism. Additionally, VRWKV-T’s RNN-like computational framework ensures a slow increase in memory usage. By the time the resolution hits $2048\times 2048$ (equivalent to 16384 tokens), VRWKV-T’s inference speed is 10 times faster than ViT-T, and its memory consumption is reduced by 80% compared to ViT-T.
We also compare the speed of our ${\rm Bi\mbox{-}WKV}$ and flash attention [7], reported in Fig. 3(b). Flash attention is highly efficient at low resolutions yet plagued by quadratic complexity, its speed decreases rapidly as the resolution increases. In high-resolution scenarios, our linear operator ${\rm Bi\mbox{-}WKV}$ demonstrates a significant speed advantage. For instance, when the input is $2048\times 2048$ (i.e., 16384 tokens) with the channel and head number settings according to ViT-B and VRWKV-B, our ${\rm Bi\mbox{-}WKV}$ operator is 2.8×2.8$\times$2.8 × faster than flash attention in inference runtime and 2.7×2.7$\times$2.7 × faster in the combined forward and backward pass.
MAE Pre-training. Similar to ViT, our VRWKV model can handle sparse inputs and benefits from the MAE pre-training [19]. By simply modifying the Q-Shift to perform a bidirectional shift operation, VRWKV can be pre-trained using MAE. The pre-trained weights can be directly fine-tuned for other tasks using a Q-Shift manner. Following the same MAE pre-training setting as ViT, and subsequent classification training similar to Sec. 4.1, our VRWKV-L achieves a top-1 accuracy improvement from 86.0% to 86.2% on ImageNet-1K val, showing the ability to acquire visual prior from masked image modeling.
@@
Figure 3: Comparison of effective receptive field (ERF) and attention runtime. (a) ERF for ViT and VRWKV in different settings. “No Shift” means no shift is used in spatial-mix and channel-mix modules. “RWKV Attn” means the original RWKV attention [35] without our modifications. Our VRWKV with Q-Shift and Bi-WKV has a more comprehensive ERF than other counterparts. (b) Attention runtime of inference (left) and forward $+$ backward (right) tested on an Nvidia A100 GPU.

### v3

@@ 4.4 Ablation Study
Settings. We conduct ablation studies of the tiny-size VRWKV on ImageNet-1K (Deng et al., 2009) to validate the effectiveness of various key components like Q-Shift and Bi-WKV. The experimental settings are consistent with Section 4.1.
@@
Table 5: Ablation on key components of the proposed VRWKV. All models are trained from scratch on ImageNet-1K. “Original” refers to the token shift in RWKV (Peng et al., 2023), which mixes tokens in a single direction.
Token Shift. We compare three approaches: no token shift, the original shift method used in RWKV, and our proposed Q-Shift. As shown in Table 5, the variation in the shift method shows performance differences. Variant 1 without token shift leads to a poor performance of 71.5, which is 3.6 points lower than our model. Even with the use of our global attention, the model using the original token shift still has a 0.7-point gap with our model.
Bidirectional Attention. The bidirectional attention mechanism enables the model to achieve global attention while the original RWKV attention has a causal mask internally. The result of Variant 3 shows the global attention mechanism brings a 2.3 points increase in the top-1 accuracy.
@@
Figure 3: Comparison of effective receptive field (ERF) and attention runtime. (a) ERF for ViT and VRWKV in different settings. “No Shift” means no shift is used in spatial-mix and channel-mix modules. “RWKV Attn” means the original RWKV attention without our modifications. Our VRWKV with Q-Shift and Bi-WKV has a more comprehensive ERF than other counterparts. (b) Attention runtime of inference (left) and forward $+$ backward (right) tested on an Nvidia A100 GPU.
Effective Receptive Field (ERF). We analyze the impact of different designs on the ERF of models based on (Ding et al., 2022) and visualize it in Figure 3(a). We visualize the ERF of the central pixel with an input size of 1024 × 1024. In Figure 3(a), “No Shift” represents the absence of the token shift method (Q-Shift), and “RWKV Attn” indicates using the original RWKV attention mechanism without our modifications for vision tasks. From the comparison in the figure, all models except the “RWKV Attn” one achieved global attention while the global capacity of the VRWKV-T model is better than that of ViT-T. Despite the assistance of Q-Shift, the central pixel in “RWKV Attn” still cannot attend to the pixels on the bottom of the image due to the large input resolution. The results of “No Shift” and Q-Shift show that the Q-Shift method expands the core range of the receptive field, enhancing the inductive bias of global attention.
Efficiency Analysis. To showcase the efficiency of our linear attention mechanism, we gradually increase the input resolution from $224\times 224$ to $2048\times 2048$ and compare the inference and memory efficiency of VRWKV-T and ViT-T. The results were tested on an Nvidia A100 GPU, as shown in Figure 1. From the curves presented in Figure 1(b), it is observed that at lower resolutions, VRWKV-T and ViT-T exhibit comparable memory usage and inference FPS. With the increase in input resolution, VRWKV-T shows a much higher speed than ViT-T. Additionally, VRWKV-T’s RNN-like computational framework ensures a slow increase in GPU memory usage. By the time the resolution hits $2048\times 2048$ (equivalent to 16384 tokens), VRWKV-T’s inference speed is 10 times faster than ViT-T, and its memory consumption is reduced by 80% compared to ViT-T. It is worth mentioning that the implementation of the Q-Shift operation in PyTorch is highly inefficient, leading to an overall decrease in model speed. However, this operation can be optimized through other means (such as using CUDA extensions). Therefore, there is still potential for further improvement in the speed of VRWKV with better engineering optimization.
We also compare the speed of our attention mechanism kernel, ${\rm Bi\mbox{-}WKV}$, with pytorch attention and flash attention (Dao et al., 2022), reported in Figure 3(b). Our Bi-WKV is significantly faster than the attention mechanism implemented using matrix multiplication (PyTorch attention), achieving a speedup of over a hundred times at a resolution of $2048\times 2048$ (i.e., 16384 tokens). Flash attention is highly optimized for memory I/O, and its matrix multiplication aligns well with the physical architecture of Nvidia GPUs, while our ${\rm Bi\mbox{-}WKV}$ lacks such hardware-level optimization. Nevertheless, in high-resolution scenarios, our ${\rm Bi\mbox{-}WKV}$ still demonstrates a significant speed advantage.
MAE Pre-training. ViTs can learn meaningful visual representations in masked image modeling (MIM). Yet, the rigor and effectiveness of linear attention vision models in this self-supervised pre-training paradigm have not been validated. Our VRWKV can handle sparse inputs and leverage MIM pre-training methods like MAE (He et al., 2017) by implementing a bidirectional shift operation that removes the vertical shift in Q-Shift. The pre-trained weights can be directly fine-tuned for other tasks using a Q-Shift manner. Following the same MAE pre-training setting as ViT, and subsequent classification training similar to Section 4.1, our VRWKV-L shows the ability to acquire visual prior from masked image modeling as shown in Table 2.
@@ 5 Conclusion
Acknowledgments
This project was supported by the National Key R&D Program of China (No. 2022ZD0161300, 2022ZD0160101), the National Natural Science Foundation of China (No. 62376134). Zhe Chen is supported by the Youth PhD Student Research Project under the National Natural Science Foundation (No. 623B2050).
Appendix A Appendix
A.1 RNN Form Forward and Backward
The attention mechanism in the spatial-mix module uses an RNN form forward and backward to achieve linear complexity of the input token number $T$. The following sections give more details of the operation.
States Recurrence Relation Initial Value
$a$ $a_{t}=\mathrm{w^{-}}\cdot a_{t-1}+e^{k_{t}}v_{t}$ $a_{-1}=0$
$b$ $b_{t}=\mathrm{w^{+}}\cdot(b_{t-1}-e^{k_{t+1}}v_{t+1})$ $b_{-1}=\sum_{i=1}^{T-1}e^{-(i-1)w+k_{i}}v_{i}$
$c$ $c_{t}=\mathrm{w^{-}}\cdot c_{t-1}+e^{k_{t}}$ $c_{-1}=0$
$d$ $d_{t}=\mathrm{w^{+}}\cdot(d_{t-1}-e^{k_{t+1}})$ $d_{-1}=\sum_{i=1}^{T-1}e^{-(i-1)w+k_{i}}$
$\frac{\mathrm{d}a}{\mathrm{d}w}$ $\frac{\mathrm{d}a_{t}}{\mathrm{d}w}=\mathrm{w^{-}}\cdot(\frac{\mathrm{d}a_{t-1}}{\mathrm{d}w}-a_{t-1})$ $\frac{\mathrm{d}a_{-1}}{\mathrm{d}w}=0$
$\frac{\mathrm{d}b}{\mathrm{d}w}$ $\frac{\mathrm{d}b_{t}}{\mathrm{d}w}=\mathrm{w^{+}}\cdot(\frac{\mathrm{d}b_{t-1}}{\mathrm{d}w}+b_{t-1}-e^{k_{t+1}}v_{t+1})$ $\frac{\mathrm{d}b_{-1}}{\mathrm{d}w}=\sum_{i=1}^{T-1}-(i-1)e^{-(i-1)w+k_{i}}v_{i}$
$\frac{\mathrm{d}c}{\mathrm{d}w}$ $\frac{\mathrm{d}c_{t}}{\mathrm{d}w}=\mathrm{w^{-}}\cdot(\frac{\mathrm{d}c_{t-1}}{\mathrm{d}w}-c_{t-1})$ $\frac{\mathrm{d}c_{-1}}{\mathrm{d}w}=0$
$\frac{\mathrm{d}d}{\mathrm{d}w}$ $\frac{\mathrm{d}d_{t}}{\mathrm{d}w}=\mathrm{w^{+}}\cdot(\frac{\mathrm{d}d_{t-1}}{\mathrm{d}w}+d_{t-1}-e^{k_{t+1}})$ $\frac{\mathrm{d}d_{-1}}{\mathrm{d}w}=\sum_{i=1}^{T-1}-(i-1)e^{-(i-1)w+k_{i}}$
$\mathrm{g}a$ $\mathrm{g}a_{t}=\mathrm{w^{+}}\cdot(\mathrm{g}a_{t-1}-\mathrm{g}y_{t}\cdot y^{\mathrm{iden}}_{t})$ $\mathrm{g}a_{0}=\sum_{i=1}^{T-1}\mathrm{g}y_{i}\cdot y_{i}^{\mathrm{iden}}\cdot e^{-(i-1)w}$
$\mathrm{g}b$ $\mathrm{g}b_{t}=\mathrm{w^{-}}\cdot\mathrm{g}b_{t-1}+\mathrm{g}y_{t-1}\cdot y^{\mathrm{iden}}_{t-1}$ $\mathrm{g}b_{0}=0$
$\mathrm{g}c$ $\mathrm{g}c_{t}=\mathrm{w^{+}}\cdot(\mathrm{g}c_{t-1}-\mathrm{g}y_{t}\cdot y^{\mathrm{iden}}_{t}\cdot y_{t})$ $\mathrm{g}c_{0}=\sum_{i=1}^{T-1}\mathrm{g}y_{i}\cdot y^{\mathrm{iden}}_{i}\cdot y_{i}\cdot e^{-(i-1)w}$
$\mathrm{g}d$ $\mathrm{g}d_{t}=\mathrm{w^{-}}\cdot\mathrm{g}d_{t-1}+\mathrm{g}y_{t-1}\cdot y^{\mathrm{iden}}_{t-1}\cdot y_{t-1}$ $\mathrm{g}d_{0}=0$
Table 6: RNN states of Bi-WKV in forward and backward process. The update in recurrence relations has a fixed number of FLOPs. $\mathrm{w}^{-}$ and $\mathrm{w}^{+}$ denotes the grow or decay vector $e^{w/T}$ and $e^{-w/T}$.The calculation of initial values is $O(TC)$ which does not affect the final complexity.
A.1.1 Backward Equation
The backward process acquires the gradient of output matrix $wkv$ (denotes as $y$) passed from the previous layer (denotes as $\mathrm{g}y$) to calculate the gradient of each input. The saved inputs are vectors $w,u\in\mathbb{R}^{C}$, key and value matrices $K,V\in\mathbb{R}^{T\times C}$ (the batch dimension is omitted). The new input is the gradient $\mathrm{g}y\in\mathbb{R}^{T\times C}$ provided by the backpropagation. The outputs include the gradients $\mathrm{g}w,\mathrm{g}u\in\mathbb{R}^{C}$, matrices $\mathrm{g}K,\mathrm{g}V\in\mathbb{R}^{T\times C}$ corresponding to the inputs, respectively. From the RNN form of the forward process, the backward can be represented in an RNN form with a linear complexity related to the token number $T$. Some intermediate variables are listed as follows:
$y_{t}^{\mathrm{num}}=a_{t-1}+b_{t-1}+e^{u+k_{t}}v_{t},\ y_{t}^{\mathrm{iden}}=1/(c_{t-1}+d_{t-1}+e^{u+k_{t}}),\ y_{t}=y_{t}^{\mathrm{num}}\cdot y_{t}^{\mathrm{iden}}.$ (10)
The outputs of backward propagation are listed as follows:
$\begin{split}\mathrm{g}w=&\sum_{t=0}^{T-1}\mathrm{g}y_{t}\cdot y_{t}^{\mathrm{iden}}(\frac{\mathrm{d}a_{t-1}}{\mathrm{d}w}+\frac{\mathrm{d}b_{t-1}}{\mathrm{d}w}-y_{t}(\frac{\mathrm{d}c_{t-1}}{\mathrm{d}w}+\frac{\mathrm{d}d_{t-1}}{\mathrm{d}w})),\end{split}$ (11)
$\begin{split}\mathrm{g}u=&\sum_{t=0}^{T-1}\mathrm{g}y_{t}\cdot y_{t}^{\mathrm{iden}}\cdot e^{u+k_{t}}(-y_{t}+v_{t}),\end{split}$ (12)
$\begin{split}\mathrm{g}k_{t}=&\mathrm{g}b_{t}\cdot e^{k_{t}}v_{t}-\mathrm{g}d_{t}\cdot e^{k_{t}}+\mathrm{g}y_{t}\cdot y_{t}^{\mathrm{iden}}(e^{k_{t}+u}v_{t}-y_{t}\cdot e^{k_{t}+u})\\ &+\mathrm{g}a_{t}\cdot e^{k_{t}}v_{t}-\mathrm{g}c_{t}\cdot e^{k_{t}},\end{split}$ (13)
$\begin{split}\mathrm{g}v_{t}=&\mathrm{g}b_{t}\cdot e^{k_{t}}+\mathrm{g}a_{t}\cdot e^{k_{t}}+\mathrm{g}y_{t}\cdot y_{t}^{\mathrm{iden}}\cdot e^{k_{t}+u}.\end{split}$ (14)
The RNN states and their initial values and recurrence relations are provided in Table 6. From the recurrence relations, all updates have a complexity of $O(C)$, which means the number of FLOPs for each update is fixed. Therefore, the final backward complexity is $O(\text{s}TC)$ where s denotes the sum of the FLOPs for all equations.
A.1.2 Implementation Details
A numerical trick to compute safe exponential in (Peng et al., 2023; Milakov & Gimelshein, 2018) is used to avoid overflow in the exponential terms of the recurrence in the forward and backward process. An example of the update of state $a$ is shown as follows:
$\begin{split}q:=\mathrm{max}(p_{t-1}-w/T,k_{t}),\end{split}$ (15)
$\begin{split}a^{\prime}=\mathrm{exp}(-w/T+p_{t-1}-q)\cdot a^{\prime}+\mathrm{exp}(k_{t}-q)\cdot v_{t},\end{split}$ (16)
$\begin{split}p_{t}=q.\end{split}$ (17)
The exponential terms in the new state $a^{\prime}$ are forced to be smaller than 1 by subtracting the max value. The subtracted part stored in $p$ is divided automatically when calculating $wkv$.
Figure 4: Performance of VRWKV and DeiT (Touvron et al., 2021a) on ImageNet-1K (Deng et al., 2009). All models are trained on a fixed resolution of $224\times 224$ and evaluated on different resolutions. Our VRWKV shows an obvious robustness on different resolutions.
Method #Param Top-1 Acc
Vim-T 7M 76.1
VRWKV-T 6M 75.1
Vim-S 26M 80.5
VRWKV-S 24M 80.1
Vim-B 98M 81.9
VRWKV-B 94M 82.0
Vim-L NA NA
VRWKV-L 335M 86.0
Table 7: Comparison with Vision Mamba (Zhu et al., 2024)(Vim) on ImageNet-1K (Deng et al., 2009). “NA” denotes not available.
A.2 Robustness on Image Resolution
Settings. In this experiment, we aim to explore whether the proposed VRWKV exhibits distinct properties compared to ViT. To this end, we evaluated the performance of different variants of DeiT (Touvron et al., 2021a) and VRWKV at different resolutions on the ImageNet-1K (Deng et al., 2009) classification task. While the training was standardized at a resolution of $224\times 224$, we evaluated the models across a range of resolutions, from $224\times 224$ to $1024\times 1024$.
Results. As shown in Figure 4, our VRWKV demonstrates stronger robustness when evaluated on a higher resolution. In contrast to DeiT (Touvron et al., 2021a), VRWKV performs better as the resolution slightly increases. For example, VRWKV-B achieved a top-1 accuracy of 82.5% at a $384\times 384$ resolution, marking an improvement of 0.5 points over its accuracy at the training resolution. When the test resolution scales up to $1024\times 1024$, VRWKV-B still maintains an accuracy of 67.2%, while DeiT-B only achieves an accuracy of 57.5%. This indicates that VRWKV has stronger potential and robustness in high-resolution scenarios and is a good alternative to ViT for high-resolution tasks.
A.3 Comparison to Vision Mamba
Figure 5: Inference time of attention mechanisms. Input resolutions are scanned from 224 to 1024. All experiments are run on Nvidia A100.
In the field of visual linear attention mechanisms, Vision Mamba (Zhu et al., 2024)(Vim) stands out as a significant implementation that has garnered considerable interest. In this study, we compare the performance and efficiency of two models in visual tasks.
Classification Performance. We compare the classification accuracy on ImageNet-1K (Deng et al., 2009). As reported in Table 7, Vim has higher Top-1 Acc in tiny and small sizes, while the base size models achieve comparable performance. Benefitting from our careful stability design, VRWKV can scale up to larger models while Vim faces instability issues in the training process.
Inference Efficiency. We compare the inference speed of three attention mechanisms: Vanilla Attn (Vaswani et al., 2017), Bi-WKV, and Vision Mamba, shown in Figure 5. As the input resolution increases, the inference cost of vanilla attention quickly surpasses that of Bi-WKV and Vision Mamba. With our optimizations and design on CUDA, our Bi-WKV demonstrates faster speeds than Vision Mamba at the same input resolution.